After an accident Nomfundo was awarded an amount from the Road Accident Fund as compensation for her injuries. She chose to receive R18 900 per month indefinitely. If money is worth 9,95% per year, compounded monthly, then the amount awarded is approximately

[1] R7 252 333.

[2] R6 565 554.

[3] R2 279 397.

[4] R189 950.

[5] none of the above


1
Expert's answer
2021-06-02T09:02:58-0400

Monthly payment of R 18900 for indefinitely@9.95% p.a

Formula for present worth"=\frac{a}{i}"

Where "I=\frac{9.95}{1200}"

"a=18900"

Present value of this payment

"=\frac{18900×1200}{9.95}"

"=R 2279397"

Therefore option 3 is the right one.